# Positivity is undecidable in tensor products of free algebras

### 09 Dec 2023 v1

@@ 4.2. Construction of $\mathcal{A}_{\mathcal{S}}(m)$
For this, suppose $N\geq n_{0}$, and once again pick an ordering on $\mathcal{X}$ and identify *$\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$, and $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ with subalgebras of $\mathcal{P}_{N}$, *$\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$, and $\mathbb{C}\mathbb{Z}_{2}^{*N}$ respectively. Let $\pi:\mathcal{P}_{N}\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ be the quotient map, and let $q=q_{2}\circ q_{1}$, where $q_{1}:\mathbb{Q}\mathbb{Z}_{2}^{*N}\to\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$ is the homomorphism sending $x_{i}\mapsto x_{i}$ for $1\leq i\leq n_{0}$, and $q_{2}:\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\to\mathcal{A}_{\mathcal{S}}(m)$ is the quotient map. Let $\widetilde{\Gamma},k^{\prime}$, and $\widetilde{W}_{m}$ be as in Proposition 6.4. For every $m\in\mathbb{N}$, define a ***-polynomial
@@
in *$\mathbb{Q}^{*}\langle\mathcal{X}\rangle\otimes_{\mathbb{Q}}\mathbb{Q}^{*}\langle\mathcal{X}\rangle\subseteq\mathcal{P}_{N}\otimes_{\mathbb{Q}}\mathcal{P}_{N}$. Then $\beta(m)$ is self-adjoint and the map $\beta:\mathbb{N}\to\mathcal{P}_{N}\otimes_{\mathbb{Q}}\mathcal{P}_{N}$ is computable by Lemmas 3.11, 4.1 and 6.4.
@@
in $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes_{\mathbb{Q}}\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\subseteq\mathbb{Q}\mathbb{Z}_{2}^{*N}\otimes_{\mathbb{Q}}\mathbb{Q}\mathbb{Z}_{2}^{*N}$. If $m\in\mathcal{S}$, then by Theorem 4.1, part (c), there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{S}}(m)$ such that $\tau(\widetilde{P}_{0}^{2})>0$. Since $\tau\circ q_{2}$ is a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ and $q_{2}(r)=0$ for all $r\in\mathcal{R}_{m}$, Lemma 6.5 implies that there is a $(0,\mathcal{X})$-synchronous state $\varphi$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that $\varphi(\widetilde{P}_{0}^{2}\otimes 1)=\tau(\widetilde{P}_{0}^{2})>0$ and $\varphi(r^{*}r\otimes 1)=\tau(q_{2}(r^{*}r))=0$ for all $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$. Then $\widetilde{\varphi}:=\varphi\circ(q_{1}\otimes q_{1})$ is a state on $\mathbb{C}\mathbb{Z}_{2}^{*N}\otimes\mathbb{C}\mathbb{Z}_{2}^{*N}$ such that $\widetilde{\varphi}\left((\pi\otimes\pi)(\beta(m))\right)=-\tfrac{1}{\widetilde{\Gamma}^{2}m^{2k^{\prime}}}\tau(\widetilde{P}_{0}^{2})<0$. We conclude that $(\pi\otimes\pi)(\beta(m))$ is not positive.
Now suppose $m\notin\mathcal{S}$. Let $\varphi$ be a state on $(\mathcal{P}_{N})_{\mathbb{C}}\otimes_{\mathbb{C}}(\mathcal{P}_{N})_{\mathbb{C}}$, and let
@@
The restriction of $\varphi$ to *$\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is an $(\epsilon,\widetilde{\mathcal{W}}_{m})$-state which is $(\epsilon,\mathcal{X})$-synchronous, so by Proposition 6.4, $\varphi(\widetilde{P}_{0}^{*}\widetilde{P}_{0}\otimes 1)=\lVert\widetilde{P}_{0}\otimes 1\rVert^{2}_{\varphi}\leq\widetilde{\Gamma}^{2}m^{2k^{\prime}}\epsilon$. Hence $\varphi(\beta(m))\geq 0$. We conclude that the complement $\overline{\mathcal{S}}$ of $\mathcal{S}$ reduces to $P_{1}(N)$. Taking $\mathcal{S}$ to be RE-hard implies that $P_{1}(N)$ is coRE-hard for $N\geq n_{0}$.
@@
The following example shows that Theorem 6.1 is not true for $\mathcal{C}=\mathbb{Z}_{2}^{*2}$.
@@
Let $G=\mathbb{Z}_{2}^{*2}$ be the infinite dihedral group. This group is amenable and linear. Hence the product group $G\times G$ is also amenable and linear. By a theorem of Bekka, $\mathbb{C}G\times G$ is residually finite-dimensional [Bek99], in the sense that an element $\alpha\in\mathbb{C}G\otimes\mathbb{C}G=\mathbb{C}G\times G$ is positive if and only if $\phi(\alpha)$ is positive for all finite-dimensional representations $\phi$. Every finite-dimensional representation of $\mathbb{C}G\times G$ decomposes into irreducible representations, so $\alpha$ is positive if and only if $\phi(\alpha)$ is positive for all irreducible finite-dimensional representations. The irreducible representations of $G$ have dimension $1$ and $2$, so the irreducible representations of $G\times G$ have dimension $1$, $2$, and $4$. Thus $\alpha\in\mathbb{C}G\times G$ is positive if and only if $\phi(\alpha)$ is positive for all representations of dimension $4$. The question of whether $\phi(\alpha)$ is not positive for all representations of dimension $4$ lies in the existential theory of the reals, and thus can be solved in $\operatorname{PSPACE}$ [Can88]. In particular, the question of whether $\alpha\in\mathbb{C}G\times G$ is positive is decidable. We do not know whether $\mathbb{C}G\times G$ is archimedean closed.

### 18 Mar 2026 v2

@@ 4.2. Construction of $\mathcal{A}_{\mathcal{L}}(m)$
For this, suppose $N\geq n_{0}$, and once again pick an ordering on $\mathcal{X}$ and identify $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, $\mathbb{C}^{*}\langle\mathcal{X}\rangle$, and $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ with subalgebras of $\mathcal{P}_{N}$, $\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$, and $\mathbb{C}\mathbb{Z}_{2}^{*N}$ respectively. Let $\pi:\mathcal{P}_{N}\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ be the quotient map, and let $q=q_{2}\circ q_{1}$, where $q_{1}:\mathbb{Q}\mathbb{Z}_{2}^{*N}\to\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$ is the homomorphism sending $x_{i}\mapsto x_{i}$ for $1\leq i\leq n_{0}$, and $q_{2}:\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\to\mathcal{A}_{\mathcal{L}}(m)$ is the quotient map. Let $\widetilde{\Gamma},k^{\prime}$, and $\widetilde{W}_{m}$ be as in Proposition 6.4. For every $m\in\mathbb{N}$, define a *-polynomial
@@
in $\mathbb{Q}^{*}\langle\mathcal{X}\rangle\otimes_{\mathbb{Q}}\mathbb{Q}^{*}\langle\mathcal{X}\rangle\subseteq\mathcal{P}_{N}\otimes_{\mathbb{Q}}\mathcal{P}_{N}$. Then $\beta(m)$ is self-adjoint and the map $\beta:\mathbb{N}\to\mathcal{P}_{N}\otimes_{\mathbb{Q}}\mathcal{P}_{N}$ is computable by Lemmas 3.11, 4.1 and 6.4.
@@
in $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes_{\mathbb{Q}}\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\subseteq\mathbb{Q}\mathbb{Z}_{2}^{*N}\otimes_{\mathbb{Q}}\mathbb{Q}\mathbb{Z}_{2}^{*N}$. If $m\in\mathcal{L}$, then by Theorem 4.1, part (c), there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ such that $\tau(\widetilde{P}_{0}^{2})>0$. Since $\tau\circ q_{2}$ is a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ and $q_{2}(r)=0$ for all $r\in\mathcal{R}_{m}$, Lemma 6.5 implies that there is a $(0,\mathcal{X})$-synchronous state $\varphi$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that $\varphi(\widetilde{P}_{0}^{2}\otimes 1)=\tau(\widetilde{P}_{0}^{2})>0$ and $\varphi(r^{*}r\otimes 1)=\tau(q_{2}(r^{*}r))=0$ for all $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$. Then $\widetilde{\varphi}:=\varphi\circ(q_{1}\otimes q_{1})$ is a state on $\mathbb{C}\mathbb{Z}_{2}^{*N}\otimes\mathbb{C}\mathbb{Z}_{2}^{*N}$ such that $\widetilde{\varphi}\left((\pi\otimes\pi)(\beta(m))\right)=-\tfrac{1}{\widetilde{\Gamma}^{2}m^{2k^{\prime}}}\tau(\widetilde{P}_{0}^{2})<0$. We conclude that $(\pi\otimes\pi)(\beta(m))$ is not positive.
Now suppose $m\notin\mathcal{L}$. Let $\varphi$ be a state on $(\mathcal{P}_{N})_{\mathbb{C}}\otimes_{\mathbb{C}}(\mathcal{P}_{N})_{\mathbb{C}}$, and let
@@
The restriction of $\varphi$ to $\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is an $(\epsilon,\widetilde{\mathcal{W}}_{m})$-state which is $(\epsilon,\mathcal{X})$-synchronous, so by Proposition 6.4, $\varphi(\widetilde{P}_{0}^{*}\widetilde{P}_{0}\otimes 1)=\lVert\widetilde{P}_{0}\otimes 1\rVert^{2}_{\varphi}\leq\widetilde{\Gamma}^{2}m^{2k^{\prime}}\epsilon$. Hence $\varphi(\beta(m))\geq 0$. We conclude that the complement $\overline{\mathcal{L}}$ of $\mathcal{L}$ reduces to $P_{1}(N)$. Taking $\mathcal{L}$ to be RE-hard implies that $P_{1}(N)$ is coRE-hard for $N\geq n_{0}$.
@@
The following example shows that Theorem 6.1 is not true for $\mathcal{C}=\mathbb{Q}\mathbb{Z}_{2}^{*2}$.
@@
Let $G=\mathbb{Z}_{2}^{*2}$ be the infinite dihedral group. This group is amenable and linear. Hence the product group $G\times G$ is also amenable and linear. By a theorem of Bekka, $\mathbb{C}G\times G$ is residually finite-dimensional [BEK99], in the sense that an element $\alpha\in\mathbb{C}G\otimes\mathbb{C}G=\mathbb{C}G\times G$ is positive if and only if $\phi(\alpha)$ is positive for all finite-dimensional representations $\phi$. Every finite-dimensional representation of $\mathbb{C}G\times G$ decomposes into irreducible representations, so $\alpha$ is positive if and only if $\phi(\alpha)$ is positive for all irreducible finite-dimensional representations. The irreducible representations of $G$ have dimension $1$ and $2$, so the irreducible representations of $G\times G$ have dimension $1$, 22, and $4$. Thus $\alpha\in\mathbb{C}G\times G$ is positive if and only if $\phi(\alpha)$ is positive for all representations of dimension $4$. The question of whether $\phi(\alpha)$ is not positive for all representations of dimension $4$ lies in the existential theory of the reals, and thus can be solved in $\operatorname{PSPACE}$ [CAN88]. In particular, the question of whether $\alpha\in\mathbb{C}G\times G$ is positive is decidable. We do not know whether $\mathbb{C}G\times G$ is archimedean closed.